OpenStudy (newarf):

what is the value of -3 * (7 + (-3)2) divided by 3

OpenStudy (psycho45):

sa -13 wckeller votre worng

OpenStudy (wckeller):

Oh wow, I messed up big time... Let me restart. (-3)(2) = -6 -6+7 = 1 -3*1=-3 -3/3=-1 Final answer: -1

OpenStudy (psycho45):

haha il y a aller

OpenStudy (newarf):

that's not a possible answer

OpenStudy (psycho45):

comment est-ce